# Neural Networks and (Virtual) Extended Formulations

### v3

@@ Abstract.
In an attempt to prove similar bounds also for general neural networks, we introduce the notion of virtual extension complexity $\operatorname{vxc}(P)$, which generalizes $\operatorname{xc}(P)$ and describes the number of inequalities needed to represent the linear optimization problem over $P$ as a difference of two linear programs. We prove that $\operatorname{vxc}(P)$ is a lower bound on the size of any neural network that optimizes over $P$. While it remains an open question to derive useful lower bounds on $\operatorname{vxc}(P)$, we argue that this quantity deserves to be studied independently from neural networks by proving that one can efficiently optimize over a polytope $P$ using a small virtual extended formulation.
@@ 1. Introduction
There is a direct translation between polytopes (as feasible sets of linear programs) and CPWL functions (represented by neural networks) through the notion of the support function $f_{P}(c)=\max_{x\in P}c^{\top}x$ of a polytope $P$. This (convex) CPWL function, which has one linear region for each vertex of $P$, uniquely determines $P$ via convex duality. Computing $f_{P}(c)$ means determining the objective value when optimizing over $P$ in $c$-direction. Each CPWL function $f_{P}$ can be represented by a neural network (Arora et al., 2018), though the required number of neurons can be large. To quantify this, we define the neural network complexity $\operatorname{nnc}(P)$ as the minimum number of neurons to represent $f_{P}$ by a maxout neural network.
@@ 1.1. Our Contributions
where $P+Q=\{p+q\mid p\in P,q\in Q\}$ is the Minkowski sum. In this definition, $P$ is a (formal) Minkowski difference of two polytopes $Q$ and $R$. Note that this is not the same as $R+(-1)\cdot Q$ but rather the inverse operation of Minkowski addition. The name virtual extension complexity is derived from virtual polytopes (Panina and Streinu, 2015), a framework for the algebraic study of formal Minkowski differences of polytopes. Observe that $\operatorname{vxc}(P)\leq\operatorname{xc}(P)$ because we can always choose $Q=\{0\}$ with $\operatorname{xc}(\{0\})=0$. In that sense, virtual extension complexity is really a strengthening of the ordinary extension complexity.
In Section 4.1, we deduce that $\operatorname{vxc}(P)$ is indeed a lower bound for $\operatorname{nnc}(P)$, up to a constant factor. This leaves the open question to find ways to lower-bound $\operatorname{vxc}(P)$ in order to achieve the original goal to lower-bound $\operatorname{nnc}(P)$. To this end, it seems to be crucial to generally obtain a better understanding of the complexity measure $\operatorname{vxc}(P)$ and its relation to $\operatorname{xc}(P)$. Observe that $P+Q=R$ is equivalent to $f_{P}=f_{R}-f_{Q}$ pointwise. Therefore, intuitively, in order to optimize over $P$, one only needs to optimize over $R$ and $Q$ and subtract the results. We make this intuition formal in Section 4.2, implying that small extended formulations for $Q$ and $R$ are sufficient to optimize efficiently over $P$. Furthermore, in Section 4.3, we provide a class of examples with $P+Q=R$ demonstrating that $\operatorname{xc}(R)$ can be much smaller than $\operatorname{xc}(P)$. This gives important insights on how extension complexity behaves under Minkowski sum and implies that we really need to look at both $\operatorname{xc}(Q)$ and $\operatorname{xc}(R)$ in order to lower-bound $\operatorname{vxc}(P)$.
@@ 1.2. Further Related Work
In this paper our lever to prove lower bounds on neural networks are combinatorial optimization problems. Complementing upper bounds were established by Hertrich and Sering (2024) for minimum spanning trees and maximum flows. Furthermore, Hertrich and Skutella (2023) proved similar upper bounds for the knapsack problem, even though they are different in flavor because some integrality assumptions are made.
@@
In terms of exact representation, it is known that a function can be represented if and only if it is CPWL (Arora et al., 2018), and it is still an open question whether constant depth is sufficient to do so (Hertrich et al., 2023; Haase et al., 2023; Grillo et al., 2025; Bakaev et al., 2025b). Interestingly, also for this question, monotone networks seem to be more amenable for proving lower bounds than their non-monotone counterparts (Valerdi, 2024; Bakaev et al., 2025a). Furthermore, the related question of how to efficiently write a non-convex CPWL function as a difference of two convex ones received quite some attention recently (Brandenburg et al., 2025; Tran and Wang, 2024).
We would like to emphasize that we view neural networks as a model of real-valued computation, as opposed to binary models of computation like Boolean circuits and Turing machines. In fact, if one restricts the inputs of a neural network to be binary, it is not too difficult to simulate AND-, OR-, and NOT-gates (Mukherjee and Basu, 2017). Thus, in such a binary model, every problem in P can be solved with polynomial-size neural networks. However, such networks would usually be very sensitive to single bits in the input. This is undesirable for practical neural networks and makes it impossible to transform these constructions naturally into exact or approximate neural networks in the real-valued model, compare the discussion by Hertrich and Sering (2024). The more useful connection to circuit complexity is through arithmetic (Shpilka et al., 2010), and in particular tropical circuits (Jukna, 2023), which are also real-valued models of computation. Again we refer to Hertrich and Sering (2024) for a more detailed discussion.
@@ 2. Preliminaries
Sometimes, a more compact way of representing a polyhedron $P$ can be obtained by representing it as the projection of a higher-dimensional polyhedron $Q\subseteq\mathbb{R}^{e}$ with $e\geq d$. Such a polytope $Q$ with $\pi(Q)=P$ for an affine projection $\pi$ is called an extended formulation of $P\subseteq\mathbb{R}^{d}$. The extension complexity of $P$ is defined as $\operatorname{xc}(P)=\min\{h(Q)\mid\pi(Q)=P\}$. The extension complexity is upper bounded by $\min\{v(P),h(P)\}$, but can be exponentially smaller. For example, the spanning tree polytope of a graph with $n$ vertices has extension complexity $\mathcal{O}(n^{3})$ (Martin, 1991), but the polytope itself has exponentially many vertices and facets in $n$. See Conforti et al. (2013) for a (not very recent) survey on extended formulations.
@@
For a polytope $P\subseteq\mathbb{R}^{d}$, its support function $f_{P}\colon\mathbb{R}^{d}\to\mathbb{R}$ is defined as $f_{P}(c)=\max_{x\in P}c^{\top}x$. In other words, it maps a linear objective direction to the optimal value obtained by optimizing over the polytope. Support functions are convex, continuous, piecewise linear, and positively homogeneous functions; and every function with these properties is a support function of a polytope. Here, a function $f\colon\mathbb{R}^{d}\to\mathbb{R}$ is positively homogeneous if $f(\lambda x)=\lambda f(x)$ for all scalars $\lambda\geq 0$. Let $\mathcal{F}^{d}$ be the set of all such functions from $\mathbb{R}^{d}\to\mathbb{R}$ and let $\mathcal{P}^{d}$ be the set of all polytopes embedded in $\mathbb{R}^{d}$. Then the map $\mathcal{P}^{d}\to\mathcal{F}^{d}$, $P\mapsto f_{P}$, is a bijection that is compatible with a certain set of operations:
@@
For an introduction to neural networks, we refer to Goodfellow et al. (2016). In this paper, we focus on so-called rank-$k$ maxout neural networks for some natural number $k\geq 2$. They generalize the well-known neural networks with rectified linear unit (ReLU) activations. Such a neural network is based on a directed, acyclic graph with node set $V$ and arcs $A\subseteq V\times V$, where the nodes are called neurons and act as computational units. The $d\geq 1$ neurons with in-degree zero are called input neurons, all other $s\geq 1$ neurons are maxout units. We assume that among the maxout units there exists exactly one neuron with out-degree zero, the output neuron. All neurons that are neither input nor output neurons are called hidden neurons. The size of the neural network is defined as the number $s$ of maxout units.
Each neuron $v$ in the neural network defines a function $z_{v}\colon\mathbb{R}^{d}\to\mathbb{R}$ as follows. Each input neuron $v$ is associated with one element $x_{i}$ of the input vector $x\in\mathbb{R}^{d}$ and simply outputs $z_{v}(x)=x_{i}$. Each maxout unit $v$ comes with a tuple of weights $w_{uv}^{i}\in\mathbb{R}$ for $i=1,\dots,k$ and all $u\in\delta^{\mathrm{in}}_{v}$, where $\delta^{\mathrm{in}}_{v}$ is the set of in-neighbors of $v$. The maxout neuron $v$ represents the following expression dependent on the outputs of its in-neighbors:
@@ 3. Lower Bounds for Monotone and Input-Convex Neural Networks
Let $\mathcal{N}_{0}$ be the maxout ICNN of size $s$ representing $f$. Without loss of generality we assume that every input neuron has a connection to every maxout unit. We can always achieve this by introducing arcs with weights equal to zero. This does not increase the number of neurons and therefore not the size of the neural network. Let $v_{1},v_{2},\dots,v_{s}$ be a topological order of the maxout units of $\mathcal{N}_{0}$.
@@
To go from $\mathcal{N}_{p-1}$ to $\mathcal{N}_{p}$ for some $p\geq 1$, we let the first p1p-1 maxout neurons unchanged. Let $I\subseteq V$ be the set of input neurons. Consider the weights $w_{uv}^{i}$ on arcs entering $v\coloneqq v_{p}$ from an input neuron $u\in I$. Let $\gamma_{u}\coloneqq\min_{i=1,\dots,k}w_{uv}^{i}$ be the smallest such weight for each $u\in I$, which might be negative or positive. Observe that
@@
Observe that the support function $f_{P}$ of a polytope $P$ is monotone if and only if $P$ is contained in the nonnegative orthant $\mathbb{R}_{\geq 0}^{d}$. Proposition 3.2 implies that for the definition of $\operatorname{mnnc}(P)$ of a polytope $P\in\mathbb{R}_{\geq 0}^{d}$, it does not matter whether we use monotone or input-convex neural networks.
We now prove Theorem 3.1 in two steps, represented by the next two propositions. The first proposition shows that small ICNNs imply small extended formulations of the epigraph. For a convex CPWL function, the epigraph is the polyhedron $\operatorname{epi}(f)=\{(x,t)\in\mathbb{R}^{d}\times\mathbb{R}\mid t\geq f(x)\}$. The proposition is related to earlier works connecting the extension complexity to Boolean and arithmetic circuit complexity, see e.g., Fiorini et al. (2021); Hrubeš and Yehudayoff (2023). A similar statement also appears in the paper introducing ICNNs, where Amos et al. (2017) prove that ICNN inference can be written as a linear program.
@@
$t\geq f(c)\ \Leftrightarrow\ t\geq\max_{x\in P}c^{\top}x\ \Leftrightarrow\ t\geq c^{\top}x\;\forall x\in P\ \Leftrightarrow\ c^{\top}x-t\leq 0\;\forall x\in P\ .$
@@
Let us analyze what we can say about the support function $f_{Q}$ of the modified polytope for $c\in\mathbb{R}_{\geq 0}^{E}$. Firstly, nonnegativity of $c$ implies $f_{Q}(c)\leq\max\{c^{\top}x\mid x\in P_{f}\cap\mathbb{R}_{\geq 0}^{E}\}\leq f(c)\leq(1+\epsilon)f_{\mathrm{MAT}}(c)$. Secondly, we claim that also $f_{Q}(c)\geq f_{\mathrm{MAT}}(c)$ holds for all $c\in\mathbb{R}_{\geq 0}^{E}$. To see this, assume the contrary, namely the existence of some $c\in\mathbb{R}_{\geq 0}^{E}$ with $f_{Q}(c)<f_{\mathrm{MAT}}(c)$. Let $x\in P_{\mathrm{MAT}}$ be an optimal vertex of the matching polytope in $c$-direction. It follows that $x\notin Q$. By definition of $Q$, it follows that $P_{f}$ and the set $X^{+}\coloneqq x+\mathbb{R}_{\geq 0}^{E}$ are disjoint. As these two sets are polyhedra, there must exist a separating hyperplane, that is, some vector $a\in\mathbb{R}^{E}$ and some value $\gamma\in\mathbb{R}$ with $a^{\top}y\leq\gamma$ for all $y\in P_{f}$ and $a^{\top}y>\gamma$ for all $y\in X^{+}$. The latter implies that $a$ is nonnegative, as the recession cone of $X^{+}$ is the nonnegative orthant. However, we then have $f_{\mathrm{MAT}}(a)\geq a^{\top}x>\gamma\geq f(a)$, contradicting the assumption of the theorem.

### v4

@@ Abstract.
In an attempt to prove similar bounds also for general neural networks, we introduce the notion of virtual extension complexity $\operatorname{vxc}(P)$, which generalizes $\operatorname{xc}(P)$ and describes the number of inequalities needed to represent the linear optimization problem over $P$ as a difference of two linear programs. We prove that $\operatorname{vxc}(P)$ is a lower bound on the size of any neural network that optimizes over $P$. While it remains an open question to derive useful lower bounds on $\operatorname{vxc}(P)$, we argue that this quantity deserves to be studied independently from neural networks by proving that one can efficiently optimize over a polytope $P$ given a virtual extended formulation with small encoding size.
@@ 1. Introduction
There is a direct translation between polytopes (as feasible sets of linear programs) and CPWL functions (represented by neural networks) through the notion of the support function $f_{P}(c)=\max_{x\in P}c^{\top}x$ of a polytope $P$, see Figure 1. This (convex) CPWL function, which has one linear region for each vertex of $P$, uniquely determines $P$ via convex duality. Computing $f_{P}(c)$ means determining the objective value when optimizing over $P$ in $c$-direction. Each CPWL function $f_{P}$ can be represented by a neural network (Arora et al., 2018), though the required number of neurons can be large. To quantify this, we define the neural network complexity $\operatorname{nnc}(P)$ as the minimum number of neurons to represent $f_{P}$ by a maxout neural network.
[…]
Figure 1. Sketch of correspondence between polytopes and support functions.
@@ 1.1. Our Contributions
where $P+Q=\{p+q\mid p\in P,q\in Q\}$ is the Minkowski sum, see Figure 2. In this definition, $P$ is a (formal) Minkowski difference of two polytopes $R$ and $Q$. Note that this is not the same as $R+(-1)\cdot Q$ but rather the inverse operation of Minkowski addition. While we start with a convex polytope $P$ in the first place, this notion of Minkowski difference does in general not result in a polytope. More precisely, it yields a polytope if and only if the normal fan of $R$ refines the normal fan of $Q$, see Postnikov et al. (2008). The name virtual extension complexity is derived from virtual polytopes (Panina and Streinu, 2015), a framework for the algebraic study of formal Minkowski differences of polytopes. Observe that $\operatorname{vxc}(P)\leq\operatorname{xc}(P)$ because we can always choose $Q=\{0\}$ with $\operatorname{xc}(\{0\})=0$. In that sense, virtual extension complexity is really a strengthening of the ordinary extension complexity.
AABBA+BA+B++==
Figure 2. Minkowski sum of two polygons.
In Section 4.1, we deduce that $\operatorname{vxc}(P)$ is indeed a lower bound for $\operatorname{nnc}(P)$, up to a constant factor. This leaves the open question to find ways to lower-bound $\operatorname{vxc}(P)$ in order to achieve the original goal to lower-bound $\operatorname{nnc}(P)$. To this end, it seems to be crucial to generally obtain a better understanding of the complexity measure $\operatorname{vxc}(P)$ and its relation to $\operatorname{xc}(P)$. Observe that $P+Q=R$ is equivalent to $f_{P}=f_{R}-f_{Q}$ pointwise. Therefore, intuitively, in order to optimize over $P$, one only needs to optimize over $R$ and $Q$ and subtract the results. We make this intuition formal in Section 4.2, implying that extended formulations for $Q$ and $R$ with small encoding size are sufficient to optimize efficiently over $P$. Furthermore, in Section 4.3, we provide a class of examples with $P+Q=R$ demonstrating that $\operatorname{xc}(R)$ can be much smaller than $\operatorname{xc}(P)$. This gives important insights on how extension complexity behaves under Minkowski sum and implies that we really need to look at both $\operatorname{xc}(Q)$ and $\operatorname{xc}(R)$ in order to lower-bound $\operatorname{vxc}(P)$.
@@ 1.2. Further Related Work
In this paper our lever to prove lower bounds on neural networks is combinatorial optimization problems. Complementing upper bounds were established by Hertrich and Sering (2024) for minimum spanning trees and maximum flows and by Hertrich et al. (2026) for regular matroids. The latter result also shows that $\operatorname{vxc}(P)\in O(n^{3})$ for $P$ being the base polytope of a regular matroid with $n$ elements. This is a first example for which $\operatorname{vxc}(P)$ is smaller than the best known upper bound on $\operatorname{xc}(P)$, which is $O(n^{6})$ in this case (Aprile and Fiorini, 2022). Furthermore, Hertrich and Skutella (2023) proved neural network constructions for the knapsack problem, even though they are different in flavor because some integrality assumptions are made.
@@
In terms of exact representation, it is known that a function can be represented if and only if it is CPWL (Arora et al., 2018), and it is still an open question whether constant depth is sufficient to do so (Hertrich et al., 2023; Haase et al., 2023; Grillo et al., 2025; Bakaev et al., 2025b). Interestingly, also for this question, monotone networks and ICNNs seem to be more amenable for proving lower bounds than their non-monotone counterparts (Valerdi, 2024; Bakaev et al., 2025a). Furthermore, the related question of how to efficiently write a non-convex CPWL function as a difference of two convex ones received quite some attention recently (Brandenburg et al., 2025; Tran and Wang, 2024). A neural network architecture that explicitly learns differences of convex functions was proposed by Sankaranarayanan and Rengaswamy (2022).
We would like to emphasize that we view neural networks as a model of real-valued computation, as opposed to binary models of computation like Boolean circuits and Turing machines. In fact, if one restricts the inputs of a neural network to be binary, it is not too difficult to simulate AND-, OR-, and NOT-gates (Mukherjee and Basu, 2017). Thus, in such a binary model, every problem in P can be solved with polynomial-size neural networks. However, such networks would usually be very sensitive to single bits in the input. This is undesirable for practical neural networks and makes it impossible to transform these constructions naturally into exact or approximate neural networks in the real-valued model, compare the discussion by Hertrich and Sering (2024). The more useful connection to circuit complexity is through arithmetic (Shpilka and Yehudayoff, 2010), and in particular tropical circuits (Jukna, 2023), which are also real-valued models of computation. For example, Jukna et al. (2020) study the effect of subtraction at different locations within tropical circuits, which is closely related to the idea of differences of monotone neural networks and ICNNs. Again, for a more detailed discussion on the relation of neural networks to arithmetic and tropical circuits, we refer to Hertrich and Sering (2024).
@@ 2. Preliminaries
Sometimes, a more compact way of representing a polyhedron $P$ can be obtained by representing it as the projection of a higher-dimensional polyhedron $Q\subseteq\mathbb{R}^{e}$ with $e\geq d$. Such a polyhedron $Q$ with $\pi(Q)=P$ for an affine projection $\pi$ is called an extended formulation of $P\subseteq\mathbb{R}^{d}$. The extension complexity of $P$ is defined as $\operatorname{xc}(P)=\min\{h(Q)\mid\pi(Q)=P\}$. The extension complexity of a polytope $P$ is upper bounded by $\min\{v(P),h(P)\}$, but can be exponentially smaller. For example, the spanning tree polytope of a graph with $n$ vertices has extension complexity $\mathcal{O}(n^{3})$ (Martin, 1991), but the polytope itself has exponentially many vertices and facets in $n$. See Conforti et al. (2013) for a (not very recent) survey on extended formulations.
@@
For a polytope $P\subseteq\mathbb{R}^{d}$, its support function $f_{P}\colon\mathbb{R}^{d}\to\mathbb{R}$ is defined as $f_{P}(c)=\max_{x\in P}c^{\top}x$, see Figure 1. In other words, it maps a linear objective direction to the optimal value obtained by optimizing over the polytope. Support functions are convex, continuous, piecewise linear, and positively homogeneous functions; and every function with these properties is a support function of a polytope. Here, a function $f\colon\mathbb{R}^{d}\to\mathbb{R}$ is positively homogeneous if $f(\lambda x)=\lambda f(x)$ for all scalars $\lambda\geq 0$. Let $\mathcal{F}^{d}$ be the set of all such functions from $\mathbb{R}^{d}\to\mathbb{R}$ and let $\mathcal{P}^{d}$ be the set of all polytopes embedded in $\mathbb{R}^{d}$. Then the map $\mathcal{P}^{d}\to\mathcal{F}^{d}$, $P\mapsto f_{P}$, is a bijection that is compatible with a certain set of operations, see Figure 4:
@@
Figure 4. The pointwise sum of support functions (blue) translates to the Minkowski sum of the associated Newton polytopes (red intervals).
@@
For an introduction to neural networks, we refer to Goodfellow et al. (2016). In this paper, we focus on so-called rank-$k$ maxout neural networks for some natural number $k\geq 2$. They generalize the well-known neural networks with rectified linear unit (ReLU) activations. Such a neural network is based on a directed, acyclic graph with node set $V$ and arcs $A\subseteq V\times V$, where the nodes are called neurons and act as computational units. The $d\geq 1$ neurons with in-degree zero are called input neurons, all other $s\geq 1$ neurons are maxout units. We assume that among the maxout units there exists exactly one neuron with out-degree zero, the output neuron. All neurons that are neither input nor output neurons are called hidden neurons. The size of the neural network is defined as the number $s$ of maxout units. We define our neural networks without biases, see Remark 2.1 below.
Each neuron $v$ in the neural network defines a function $z_{v}\colon\mathbb{R}^{d}\to\mathbb{R}$ as follows. Each of the $d$ input neurons is associated with exactly one element $x_{i}$ of the input vector $x\in\mathbb{R}^{d}$. Such an input neuron $v$ simply outputs $z_{v}(x)=x_{i}$. Each maxout unit $v$ comes with a tuple of weights $w_{uv}^{i}\in\mathbb{R}$ for $i=1,\dots,k$ and all $u\in\delta^{\mathrm{in}}_{v}$, where $\delta^{\mathrm{in}}_{v}$ is the set of in-neighbors of $v$. The maxout neuron $v$ represents the following expression dependent on the outputs of its in-neighbors:
@@ 3. Lower Bounds for Monotone and Input-Convex Neural Networks
Let $\mathcal{N}_{0}$ be the maxout ICNN of size $s$ representing $f$. Without loss of generality we assume that every input neuron has a connection to every maxout unit. We can always achieve this by introducing arcs with weights equal to zero. This does not increase the number of neurons and therefore not the size of the neural network according to our definition of size. [^3] However, it does potentially increase the number of arcs and therefore (after modifying the weights) the number of parameters of the network. Let $v_{1},v_{2},\dots,v_{s}$ be a topological order of the maxout units of $\mathcal{N}_{0}$.
@@
To go from $\mathcal{N}_{p-1}$ to $\mathcal{N}_{p}$ for some $p\geq 1$, we leave the first p1p-1 maxout neurons unchanged. Let $I\subseteq V$ be the set of input neurons. Consider the weights $w_{uv}^{i}$ on arcs entering $v\coloneqq v_{p}$ from an input neuron $u\in I$. Let $\gamma_{u}\coloneqq\min_{i=1,\dots,k}w_{uv}^{i}$ be the smallest such weight for each $u\in I$, which might be negative or positive. Observe that
@@
Observe that the support function $f_{P}$ of a polytope $P$ is monotone if and only if $P$ is contained in the nonnegative orthant $\mathbb{R}_{\geq 0}^{d}$. Proposition 3.2 implies that for the definition of $\operatorname{mnnc}(P)$ of a polytope $P\subseteq\mathbb{R}_{\geq 0}^{d}$, it does not matter whether we use monotone or input-convex neural networks.
We now prove Theorem 3.1 in two steps, represented by the next two propositions. The first proposition shows that small ICNNs imply small extended formulations of the epigraph. For a convex CPWL function, the epigraph is the polyhedron $\operatorname{epi}(f)=\{(x,t)\in\mathbb{R}^{d}\times\mathbb{R}\mid t\geq f(x)\}$. The proposition is related to earlier works connecting the extension complexity to Boolean and arithmetic circuit complexity, see e.g., Fiorini et al. (2021); Hrubeš and Yehudayoff (2023), and also the result by Goemans (2015) on using sorting networks for constructing an extended formulation of the permutahedron. A similar statement also appears in the paper introducing ICNNs, where Amos et al. (2017) prove that ICNN inference can be written as a linear program.
@@
$t\geq f_{P}(c)\ \Leftrightarrow\ t\geq\max_{x\in P}c^{\top}x\ \Leftrightarrow\ t\geq c^{\top}x\;\forall x\in P\ \Leftrightarrow\ c^{\top}x-t\leq 0\;\forall x\in P\ .$
@@
Let us analyze what we can say about the support function $f_{Q}$ of the modified polytope for $c\in\mathbb{R}_{\geq 0}^{E}$. Firstly, nonnegativity of $c$ implies $f_{Q}(c)\leq\max\{c^{\top}x\mid x\in P_{f}\cap\mathbb{R}_{\geq 0}^{E}\}\leq f(c)\leq(1+\epsilon)f_{\mathrm{MAT}}(c)$. Secondly, we claim that also $f_{Q}(c)\geq f_{\mathrm{MAT}}(c)$ holds for all $c\in\mathbb{R}_{\geq 0}^{E}$. To see this, assume the contrary, namely the existence of some $c\in\mathbb{R}_{\geq 0}^{E}$ with $f_{Q}(c)<f_{\mathrm{MAT}}(c)$. Let $x\in P_{\mathrm{MAT}}$ be an optimal vertex of the matching polytope in $c$-direction. It follows that $x\notin Q$. As $P_{\mathrm{MAT}}\subseteq\mathbb{R}_{\geq 0}^{E}$, by definition of $Q$, it follows that $P_{f}$ and the set $X^{+}\coloneqq x+\mathbb{R}_{\geq 0}^{E}$ are disjoint. As these two sets are polyhedra, there must exist a separating hyperplane, that is, some vector $a\in\mathbb{R}^{E}$ and some value $\gamma\in\mathbb{R}$ with $a^{\top}y\leq\gamma$ for all $y\in P_{f}$ and $a^{\top}y>\gamma$ for all $y\in X^{+}$. The latter implies that $a$ is nonnegative, as the recession cone of $X^{+}$ is the nonnegative orthant. However, we then have $f_{\mathrm{MAT}}(a)\geq a^{\top}x>\gamma\geq f(a)$, contradicting the assumption of the theorem.